HOW DO YOU FIND THE ANSWER TO 52x+33x+18+2=100x+1+3
Answer by checkley77(12844)

52x+33x+18+2=100x+1+3
85x-100x=4-20
-15x=-16
x=16/15 answer.
Proof:
52(16/15)+33(16/15)+20=100(16/15)+4
(52+33-100)16/15=-16
-15(16/15)=-16
-16=-16